OpenStudy (anonymous):

ill give a medal:) Using complete sentences, describe to the Martians how to find the inverse of your function. My function is f(x)=4x+2

OpenStudy (anonymous):

I hated the hell out of this assignment Unless you look through the lessons they half explain, you'll really never understand this class ._. It gets a lot harder from this point, so pay attention before you start slacking behind because you can't find the answers online.

OpenStudy (studygurl14):

f(x) = 4x + 2 y = 4x + 2 Switch the x and y values x = 4y + 2 Solve for y x - 2 = 4y \[y=\frac{ x-2 }{4 }\]

OpenStudy (studygurl14):

So... \[f^{-1}(x)=\frac{x-2}{4}\]

OpenStudy (anonymous):

lol that's what everyone has been telling me. and my teacher is a feather too-_-

OpenStudy (anonymous):

I'm only about 31% done and I'm in week 7. It takes less time to try to learn what they're teaching than to look online. It benefits you a lot when you go through the lesson for the DBA's. Open Study helps and I really appreciate that StudyGurl didn't just hand you the answer or lecture you